OpenStudy (anonymous):

find the slope and y intercept for the given equation: x + y = 6 m= b= find the slope and y intercept for the given equation: x + y = 6 m= b= @Mathematics

OpenStudy (anonymous):

m = -1 b = 6

OpenStudy (anonymous):

How did you get this?

OpenStudy (anonymous):

y = -1x + 6

OpenStudy (anonymous):

Ok thanks.

OpenStudy (anonymous):

rearrange to the form y = mx + c

OpenStudy (anonymous):

tw

OpenStudy (anonymous):

2x + 5y = 8

OpenStudy (anonymous):

5y = -2x + 8 2 8 y = - --- x + ---- 5 5 2 8 so slope is - ---- and y intercept is ---- 5 5